OpenStudy (anonymous):

The function defined by f(x)=x^3-3x^2 for all real numbers x has a relative maximum at x= ?

OpenStudy (turingtest):

max and mins occur at f'(x)=0 f(x)=x^3-3x^2 f'(x)=3x^2-6x=3x(x-2)=0--->x={0,2} to find if this is a min or max we must look at f''(x) f''(x)=6x-6=0--->x=1 so for x<1 f''(x)<0 (concave down) and for x>1 f''(x)>0 (concave up) since the critical point x=0<1is in a concave down region it is a relative max.